Airline Patton travels 135.45 miles in 15 minutes. airline Galperin travels 176 miles in 20 minutes. How many more miles does Patton travel than Galperin in 1 hour?
Answer by richwmiller(17219)

p=4*135.45,
g=3*176,
m=p-g
g=528, p=541.8
m=13.8 miles more,